OpenStudy (anonymous):

Find the number of days it takes for 1/2 of the substance to remain.1/4 to remain. 12.5 to remain

OpenStudy (aaronq):

you need some more data. A "half-life" is an amount of time, it is the amount of time where half of the original amount was reduced by half. For example, suppose you have 100 g of a substance with a half-life of 1 hour. At the beginning (t=0), you have 100g After 1 hours (t=1 h) you have: \(\dfrac{100~g}{2}=50~ g\) After 2 hours (t=2 h) you have: \(\dfrac{50~g}{2}=25~ g\) After 3 hours (t=3 h) you have: \(\dfrac{25~g}{2}=12.5~ g\)